• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Mindy

    Log entry time 08:01:08 on June 9,2010

    Entry number 325380

    keyword=06/09/10 shift summary

    Due to the IOC problem, we only took the left-arm data.
    We requested MCC for the harp scan around 5:30 AM.

    Before the harp scan:
    =====================
    prod.
    4417 OK multiplets:78K; 128.1mC
    4418 OK multiplets:88K; 145.7mC
    4419 OK multiplets:88.2K; 145.7mC
    4420 OK multiplets:78.6K; 129.8mC
    4423 OK multiplets:88.8K; 132.2mC
    4424 OK multiplets:69.0K; 114.1mC
    4425 OK multiplets:79.1K; 132.2mC

    after the harp scan:
    ====================
    1) move the target to the empty
    2) beam = 5uA
    3) turn the septum off
    4) turn the beam off ==> for the reduction of radiation level
    5) turn the right-arm dipole off arounf 7:30 AM
    ==> for the reduction of magnetic field

    today's run plan:
    1) ask for the control access to fix IOC/right-arm dipole problem
    may take 4 - 6 hours
    2) tune Compton chicane; figure out why Compton background rate is
    high